OpenStudy (anonymous):

WILL GIVE MEDAL!!! Rewrite the equation below in standard form. 2/3x=-5/2y+2

OpenStudy (ankit042):

y = mx+c now you can rearrange your equation.

OpenStudy (ankit042):

@cynthia_franco you have any questions?

OpenStudy (anonymous):

standard form of a line is Ax+By=C

OpenStudy (anonymous):

A is a positive integer and B and C are integers

OpenStudy (anonymous):

15x+4y=-4

OpenStudy (anonymous):

@ankit042 y=mx+c is not called the standard form..its just another way of expressing the equation..:D

OpenStudy (ankit042):

@Krishnadas My bad I didn't see the question properly.

OpenStudy (anonymous):

:D